# CoSQA+: Enhancing Code Search Evaluation with a Multi-Choice Benchmark and Test-Driven Agents

### v5

@@ V-B RQ2: Performance of Agent-based Annotator
To establish a reliable ground truth, we first have three Python experts design test programs for 400 randomly selected query-code pairs and annotate them based on test execution results. The final ground truth labels are determined by majority voting among these three experts. We then evaluate three different annotation approaches against this ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet).
For measuring the agreement across three Python experts with test programs, we use Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators [61].
@@
Agent-based Annotator 0.920
Human Wo. Test Program 0.880
Deepseek-V3 0.865
O1-mini 0.833
Claude-3.5-Sonnet 0.845
As shown in Table IV, our agent-based annotator achieves 92.0% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.0%), Deepseek-V3 (84.0%), O1-mini (82.0%) and Claude-3.5-Sonnet (84.0%). This high accuracy demonstrates that our automated agent-based approach can effectively match the quality of expert annotations with test program verification.
@@
In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.818, notably higher than human annotations without test program verification (0.61), which suggests that test programs help reduce subjectivity in the annotation process.
RQ2 Summary: Our agent-based annotator achieves 92.0% accuracy, outperforming both traditional human annotation and single LLM approaches. The high executability rate (83.67%) of generated test programs demonstrates the reliability of our automated annotation approach.
@@ V-C RQ3: Existing Methods Performance on CoSQA+
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark using NDCG@10 as the primary metric. NDCG@10 is particularly suitable for our scenario, where multiple correct code snippets may exist for a query, as it measures how well a model ranks all relevant results rather than focusing only on the first correct match. Additionally, we report Mean Reciprocal Rank (MRR) to assess the ability to rank the first relevant result highly. To further comprehensively evaluate the performance of the models, we also introduce Mean Average Precision (MAP) and Recall as additional metrics. MAP provides a measure of the overall ranking quality by averaging the precision scores across all relevant results, which is especially useful for scenarios where the order of multiple correct matches matters. Meanwhile, Recall evaluates the model’s ability to retrieve all relevant code snippets from the dataset, ensuring that no important results are missed. Together, these metrics offer a more robust assessment of the model’s effectiveness in semantic code search tasks. The tested methods include both traditional information retrieval techniques and advanced deep learning-based models. Specifically, we compare a traditional Bag of Words (BoW) approach [27] with state-of-the-art semantic code search models, namely: CodeBERT [12], UniXcoder [31], and CodeT5+ embedding [34]. In addition, we evaluate several recent sentence embedding models, including Jina-embedding-v3 [35], All-MiniLM-L12-v2 [40], All-mpnet-base-v2 [41], and Multilingual-e5-large [38].
@@
From the evaluation results, the performance differences among embedding models seem not significant. However, due to the richness of CoSQA+, which effectively eliminates random factors, even subtle differences in metrics can reflect the overall capability gap between models. Moreover, embedding models generally perform poorly on our benchmark. From the perspective of model design, embedding models are mainly used for retrieval tasks. To improve inference speed, their weight parameters are usually small, limiting their ability to deeply understand query sentences and code.
@@
RQ3 Summary: The evaluation on CoSQA+ clearly indicates that deep learning-based models significantly outperform traditional retrieval methods for semantic code search. Notably, CodeBERT leads the performance ranking by achieving the highest scores in both NDCG@10 and MRR, while the alternative embedding-based models also offer competitive results.
@@ V-D RQ4: Contributions of Different Components in Test-Driven Agents
Single LLM 0.865
PS + FA 0.840
PS + TPG + FA 0.833
PS + TPG + TE + FA 0.835
PS + TPG + Simulated Executor + FA 0.840
PS + TPG + TE + BF + FA 0.920
@@
The results indicate that the baseline Single LLM achieves an accuracy of 0.865. When the system uses only the Preliminary Screener in combination with the Final Arbiter, the accuracy slightly drops to 0.840. Incorporating the Test Program Generator alongside these yields an accuracy of 0.833, suggesting that generating test programs without execution may introduce noise or irrelevant information, which can harm performance [62]. Adding the Test Executor to the configuration improves the performance marginally (0.835), and using a Simulated Test Executor does not provide a significant advantage (0.840). This aligns with the observation that simulated execution cannot fully capture the complexity of the real compile-and-run process, potentially introducing inaccuracies or misleading feedback. Notably, the inclusion of the Bug Fixer increases the accuracy substantially to 0.920. This enhancement demonstrates that the Bug Fixer is crucial for processing the test programs effectively, as it enables the system to generate real execution results that are vital for improving prediction accuracy.
@@ VI-A Test Misjudgment
Previous studies show that as the size of the codebase increases, the challenges faced by embedding models also increase [63, 64]. This conclusion aligns with intuition, but with the analysis of other datasets and the construction of CoSQA+, we find that the decline in performance is not only due to the need to exclude more confusing codes but also because of the increased likelihood of test misjudgments. For example, a LLM may select a piece of code that meets the query requirements, but this query-code pair is not included in the candidate code set. As a result, this correct selection is misjudged as incorrect.
@@ VII Conclusion
In this paper, we present CoSQA+, the first large-scale multi-choice code search dataset with 412,080 pairs. Our dataset expands from single-choice to multi-choice query-code matching pairs, enhancing its suitability for real-world code search tasks. Through rigorous experiments, CoSQA+ demonstrates superior quality and training performance compared to existing benchmarks like CoSQA. The test-driven annotation approach achieves a high accuracy of 92.0%, outperforming traditional human annotation and single LLM methods. Experiments on CoSQA+ show that CodeBERT is better suited for multi-choice code search tasks. We believe our agents-annotated CoSQA+ benchmark will offer insights for other research that involves natural language and code, and for evaluations on a multi-choice code search task.

### v6

@@ V-B RQ2: Performance of Agent-based Annotator
To establish a reliable ground truth, we recruit three experienced Python developers with 3-5 years of professional programming experience and computer science backgrounds to serve as expert annotators. These experts first independently design test programs for 1000 randomly selected query-code pairs from CoSQA+, following a standardized evaluation protocol that required: (1) Functional correctness verification through executable test cases. (2) Edge case coverage assessment. (3) Query requirement satisfaction analysis. They then annotate the selected pairs based on test execution results.
The final ground truth labels are determined by majority voting among these three experts, with an initial inter-annotator agreement measured by Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators [63]. In cases of disagreement, experts engage in discussion sessions to reach a consensus. We then evaluate three different annotation approaches against this established ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet), each provided with identical query-code pairs and evaluation criteria.
@@
Agent-based Annotator 0.939±0.020
Human Wo. Test Program 0.891±0.026
Deepseek-V3 0.883±0.027
O1-mini 0.881±0.027
Claude-3.5-Sonnet 0.863±0.029
As shown in Table IV, our agent-based annotator achieves 93.9% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.1%), Deepseek-V3 (88.3%), O1-mini (88.1%) and Claude-3.5-Sonnet (86.3%). This high accuracy demonstrates that our automated agent-based approach can effectively match the quality of expert annotations with test program verification. To ensure appropriate interval estimation for our data characteristics, we adopt the Clopper–Pearson exact method with 95% confidence intervals for accuracy metrics in TABLE IV. Clopper–Pearson exact method is well-suited for moderate sample sizes and high sample proportions (i.e., proportions close to 0 or 1).
@@
In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.82, notably higher than human annotations without test program verification (0.60), which suggests that test programs help reduce subjectivity in the annotation process.
RQ2 Summary: Our agent-based annotator achieves 93.9% accuracy, outperforming both traditional human annotation and single LLM approaches. The high executability rate (83.67%) of generated test programs demonstrates the reliability of our automated annotation approach.
@@ V-C RQ3: Existing Methods Performance on CoSQA+
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark using MAP@10 as the primary metric. MAP@10 is particularly suitable for our scenario, where multiple correct code snippets may exist for a query, as it measures how well a model ranks all relevant results rather than focusing only on the first correct match. Additionally, we report Mean Reciprocal Rank (MRR) to assess the ability to rank the first relevant result highly.
To provide a more comprehensive evaluation, we also include NDCG [64, 22, 65, 66] (Normalized Discounted Cumulative Gain) and Recall as additional metrics. Although NDCG is typically used for graded relevance judgments, in our case where relevance is binary (either matching or not), it still serves as a useful metric by emphasizing the importance of ranking relevant results higher in the list. Meanwhile, Recall measures the model’s ability to retrieve all relevant code snippets from the dataset, ensuring comprehensive coverage of correct results. Together, these metrics offer a robust and multi-faceted assessment of the model’s performance in semantic code search tasks. The tested methods include both traditional information retrieval techniques and advanced deep learning-based models. Specifically, we compare a traditional Bag of Words (BoW) approach [27] with state-of-the-art semantic code search models, namely: CodeBERT [12], UniXcoder [31], and CodeT5+ embedding [34]. In addition, we evaluate several recent sentence embedding models, including Jina-embedding-v3 [35], All-MiniLM-L12-v2 [40], All-mpnet-base-v2 [41], and Multilingual-e5-large [38].
@@
Regarding the evaluation metrics (MRR, NDCG, MAP): these inherently assign high weight to the top-ranked results. Consequently, the most critical factor influencing a model’s score is its ability to rank correct candidate snippets at the very top positions. The internal relative order of matched code snippets in the search results has no influence on the score. A model is considered better on CoSQA+ if it can find more matched code and position them at the top of the results.
From the evaluation results, the performance differences among embedding models do not seem significant. However, due to the richness of the CoSQA+, which effectively eliminates random factors, even subtle differences in metrics can reflect the overall capability gap between models. Moreover, embedding models generally perform poorly on our benchmark. From the perspective of model design, embedding models are mainly used for retrieval tasks. To improve inference speed, their weight parameters are usually small, limiting their ability to deeply understand query sentences and code.
@@
RQ3 Summary: The evaluation on CoSQA+ clearly indicates that deep learning-based models significantly outperform traditional retrieval methods for semantic code search. Notably, CodeBERT leads the performance ranking by achieving the highest scores in both MAP@10 and MRR, while the alternative embedding-based models also offer competitive results.
@@ V-D RQ4: Contributions of Different Components in Test-Driven Agents
Single LLM 0.883
PS + FA 0.873
PS + TPG + FA 0.852
PS + TPG + TE + FA 0.854
PS + TPG + Simulated Executor + FA 0.853
PS + TPG + TE + BF + FA 0.939
@@
The results indicate that the baseline Single LLM achieves an accuracy of 0.883. When the system uses only the Preliminary Screener in combination with the Final Arbiter, the accuracy slightly drops to 0.873. Incorporating the Test Program Generator alongside these yields an accuracy of 0.852, suggesting that generating test programs without execution may introduce noise or irrelevant information, which can harm performance [67]. Adding the Test Executor to the configuration improves the performance marginally (0.854), and using a Simulated Test Executor does not provide a significant advantage (0.853). This aligns with the observation that simulated execution cannot fully capture the complexity of the real compile-and-run process, potentially introducing inaccuracies or misleading feedback. Notably, the inclusion of the Bug Fixer increases the accuracy substantially to 0.939. This enhancement demonstrates that the Bug Fixer is crucial for processing the test programs effectively, as it enables the system to generate real execution results that are vital for improving prediction accuracy.
@@
V-E RQ5: Cross-Language Generalization of the Test-Driven Agent
This experiment is to explore the generalizability of the test-driven agent’s annotation performance beyond Python.
Experimental Setup
We collect and annotate query-code pairs from CodeSearchNet (CSN) for PHP, Java, and Go to benchmark the test-driven and single-LLM annotation methods. We select these languages because (1) they rank among the top ten most-used languages, ensuring real-world relevance [68], and (2) they represent distinct paradigms: dynamically typed scripting (PHP), statically typed object-oriented (Java), and statically typed procedural/concurrent (Go). We build the benchmark in two steps. First, we extract unannotated query-code pairs from CSN for each language. Second, we label every pair using the same pipeline described in RQ2 to ensure consistent evaluation. To evaluate cross-language performance, we adapt the pipeline in Algorithm 1 to each language. We slightly refine the prompts and configure separate execution environments for PHP, Java, and Go. We then compare two annotation methods against the human-labeled ground truth: (1) the test-driven agent based on DeepSeek-V3 and (2) the single-pass LLM annotator (DeepSeek-V3). Accuracy is the metric, defined as the fraction of labels that match human annotations.
Result and Analysis
The results reveal a consistent pattern: the test-driven agent achieves higher annotation accuracy than single LLM annotators across all three non-Python languages, with significant improvements observed in PHP (82.89% vs. 72.70%, +12.17%), Java (79.96% vs. 69.59%, +10.37%), and Go (81.93% vs. 73.49%, +8.44%). This performance improvement indicates that the test-driven framework, by incorporating functional validation and error correction mechanisms, effectively enhances cross-language code understanding capabilities, outperforming the baseline of single LLMs.
Our test-driven agent consistently outperforms the single LLM baseline in accuracy. This highlights a key advantage: our approach mitigates the limitations of pre-trained models like DeepSeek-V3. Such models may exhibit inherent limitations stemming from their training data composition, which often over-represents certain languages or paradigms. This can reduce their robustness when handling syntax and features specific to other languages. For instance, PHP’s dynamic typing quirks pose interpretive challenges for the single LLM baseline. However, the test-driven agent’s structured validation process effectively addresses the issues, resulting in clearer improvements. Notably, PHP shows the largest accuracy gain (12.17%) compared to the baseline, which could stem from its unique blend of scripting and web-focused features—elements that demand stricter contextual validation.
RQ5 Summary: The test-driven agent outperforms single LLM annotators across PHP, Java, and Go with an average 10.33% accuracy gain, showing strong cross-language generalization.
@@ VI-A Test Misjudgment
Previous studies show that as the size of the codebase increases, the challenges faced by embedding models also increase [69, 70]. This conclusion aligns with intuition, but with the analysis of other datasets and the construction of CoSQA+, we find that the decline in performance is not only due to the need to exclude more confusing codes but also because of the increased likelihood of test misjudgments. For example, an LLM may select a piece of code that meets the query requirements, but this query-code pair is not included in the candidate code set. As a result, this correct selection is misjudged as incorrect.
@@ VI-C Ethical Consideration
VI-D Failure Case Analysis
Our test-driven agents achieve 93.9% accuracy, but understanding the remaining 6.1% failure cases is crucial for future improvements. We analyze ten false-positive and false-negative cases and identify four primary root causes: (1) Ambiguous queries (30% of cases) where natural language descriptions like "python fabric verbose logging" have multiple valid interpretations (e.g., generic logging functionality vs. Fabric library-specific SSH logging), leading to mismatched test generation; (2) Human annotation biases (30% of cases) where annotator subjectivity leads to incorrect judgments, such as being misled by superficial keyword matching while agents correctly identify functional mismatches; (3) Semantic misalignment (20% of cases), where agents fail to capture the precise intent behind technical terms or behavioral descriptors; (4) Validation misalignment (20% of cases), where test programs validate code functionality rather than query requirements, such as testing all features of a function instead of the specific query need. The analysis reveals that most failure cases stem from inherent limitations in natural language understanding rather than systematic flaws in our test-driven approach. Query ambiguity remains the most significant challenge, suggesting that future work should focus on multi-interpretation test generation or query clarification mechanisms. Importantly, our agents sometimes perform better than ground truth labels, indicating that test-driven verification can identify annotation failure cases in human-labeled datasets.
VI-E Validation Strategy and Scalability Constraints
Complete human verification of large-scale datasets faces fundamental scalability barriers. Expert annotation with test programs requires 34 minutes per pair, making exhaustive validation of 412,080 pairs economically prohibitive at approximately 234,000 person-hours. Our sampling-based validation strategy addresses this constraint while maintaining methodological rigor. To strengthen validation beyond our 1,000-sample ground truth evaluation, we implemented an additional verification mechanism: a specialized checker agent using DeepSeek-V3 that performs secondary validation of our test-driven annotations. This checker evaluates the complete annotation pipeline output, including preliminary screening results, test programs, execution outcomes, and final judgments. The checker achieves high agreement with human expert annotations (Krippendorff’s $\alpha$ = 0.914) and, as ground truth, grants the test-driven agents 92.8% accuracy. The high inter-annotator agreement between our checker and human experts demonstrates that automated verification can reliably assess annotation quality at scale. This checker-based validation mechanism enables quality assurance across the entire dataset while maintaining economic feasibility, offering a practical solution to the scalability constraints of traditional manual verification approaches.
@@ VII Conclusion
In this paper, we present CoSQA+, the first large-scale multi-choice code search dataset with 412,080 pairs. Our dataset expands from single-choice to multi-choice query-code matching pairs, enhancing its suitability for real-world code search tasks. Through rigorous experiments, CoSQA+ demonstrates superior quality and training performance compared to existing benchmarks like CoSQA. The test-driven annotation approach achieves a high accuracy of 93.9%, outperforming not only traditional human annotation and single-LLM methods but also demonstrating strong cross-language generalizability. Experiments on CoSQA+ show that CodeBERT is better suited for multi-choice code search tasks. We believe our agents-annotated CoSQA+ benchmark will offer insights for other research that involves natural language and code, and for evaluations on a multi-choice code search task.